# ReSplat: Learning Recurrent Gaussian Splats

### v1

@@ 4.1 Main Results
Table 1: Evaluation of 8 input views ($512\times 960$) on DL3DV. The standard optimization-based approach 3DGS (Kerbl et al., 2023) requires several thousands of iterations to reach convergence, while our feed-forward ReSplat is significantly faster and is able to benefit from additional iterations. Previous per-pixel feed-forward models MVSplat (Chen et al., 2024b) and DepthSplat (Xu et al., 2025) produces millions of Gaussians, while our ReSplat compresses the number of Gaussians by $16\times$, also leading to $4\times$ faster rendering speed. Time is measured on a single GH200 GPU.
@@
1 26.69 0.849 0.176 246K 0.440 0.0007
2 26.80 0.850 0.174 246K 0.569 0.0007
3 26.82 0.850 0.174 246K 0.693 0.0007
Figure 3: Visual comparisons on DL3DV. Our ReSplat renders higher quality images than both optimization and feed-forward methods.
@@
8 Views at $512\times 960$ Resolution on DL3DV. We report the results on the DL3DV benchmark split (140 scenes) in Table˜1. For experiments with 3DGS (Kerbl et al., 2023), we perform per-scene optimization on the 8 input views for all the 140 scenes, while for feed-forward models, we perform zero-shot feed-forward inference on the unseen scenes after training on the DL3DV training set. We observe that 3DGS optimization typically converges with 4K optimization steps, and optimizing longer could lead to overfitting due to the sparse input views, thus we report the best results at 4K iterations. As shown in Table˜1, 3DGS optimization is considerably slow due to the large number of iterations required, while our feed-forward ReSplat is $100\times$ faster and is able to benefit from additional iterations. We observe that our recurrent model converges after 3 iterations, which also makes it efficient thanks to the fast convergence. In addition, previous per-pixel feed-forward models MVSplat (Chen et al., 2024b) and DepthSplat (Xu et al., 2025) produces millions of Gaussians, while our ReSplat compresses the number of Gaussians by $16\times$, also leading to $4\times$ faster rendering speed. Overall, our ReSplat outperforms 3DGS by 3.3dB PSNR and DepthSplat by 2.6dB PSNR with superior efficiency on the number of Gaussians and the rendering speed. The visual comparisons are shown in Fig.˜3, which demonstrates the higher rendering quality of our method. We provide additional visualizations of reconstructed Gaussians in appendix (Fig.˜7 and Fig.˜8).
Generalization Across Datasets and Image Resolutions. We further evaluate the generalization abilities of our feed-forward model trained on DL3DV at $512\times 960$ resolution. First, when generalizing to the unseen RealEstate10K dataset (Fig.˜4(a)), the improvement brought by our recurrent model is more significant since unlike single-step feed-forward models (iteration 0), our model is able to adapt to the test data using the rendering error and thus reduces the domain gap. Second, existing single-step feed-forward models usually suffer from performance drop when the testing image resolution is different from training. However, with our recurrent model, we can significantly improve the robustness to different testing resolutions (Fig.˜4(b)). For example, our recurrent model improves 4dB PSNR when generalizing from $512\times 960$ to $320\times 640$. Both experiments demonstrate the benefits of our recurrent model, which is able to adapt to the unseen test scenario using the rendering error as a feedback signal and thus greatly improves the robustness.
(a) Cross-dataset generalization. Our DL3DV-trained recurrent model improves more when generalizing to the unseen RealEstate10K dataset.
(b) Cross-resolution generalization. Our recurrent model is more robust to unseen image resolutions that are different from training ($512\times 960$).
@@
Figure 4: Generalization to unseen datasets and image resolutions. The improvement brought by our recurrent model is more pronounced when generalizing to unseen scenarios. The model is trained on DL3DV at $512\times 960$ resolution. We note that the RealEstate10K dataset is easier than DL3DV.
@@
Table 2: Evaluation of 16 input views ($540\times 960$) on DL3DV. ReSplat produces $4\times$ fewer Gaussians than Long-LRM but still outperforms it. The results of other baselines are borrowed from Long-LRM.
@@
ReSplat (Ours) 0 22.69 0.742 0.307 0.7sec 518K
1 23.12 0.752 0.296 1.2sec 518K
2 23.23 0.755 0.293 1.7sec 518K
@@
ReSplat (Ours) 29.72 0.911 0.100
Table 3: Evaluation of two input views ($256\times 256$) on RealEstate10K. ReSplat outperforms all previous feed-forward 3DGS models. Compared to the 3DGS-free model LVSM, results are similar, but ReSplat offers $20\times$ faster rendering speed thanks to the 3DGS representation.
@@
16 Views at $540\times 960$ Resolution on DL3DV. We follow Long-LRM (Chen et al., 2025b) for this evaluation setup such that a direct comparison is possible. The results of 3DGS (Kerbl et al., 2023), Mip-Splatting (Yu et al., 2024), and Scaffold-GS (Lu et al., 2024) are borrowed from Long-LRM paper. This experiment aims to reconstruct the full DL3DV scene from 16 input views, which is particularly challenging due to the large scale of the DL3DV dataset. However, our ReSplat still outperforms previous optimization and feed-forward methods, as shown in Table˜2. Notably, Long-LRM uses Gaussian pruning based on the opacity values during training and evaluation, which leads to $4\times$ reduction of number of Gaussians. In contract, we compress the Gaussians by $16\times$, thus our final reconstruction has $4\times$ fewer Gaussians than Long-LRM while still outperforming it. Our reconstruction time is slower than Long-LRM mainly because of the kNN operation, which is expensive to compute with large number (e.g., >500>500K) of points. Further optimizing the implementation could potentially improve our reconstruction speed.
2 Views at $256\times 256$ Resolution on RealEstate10K. Considering the redundancy is not a major issue with two input views at low resolution ($256\times 256$), we perform $4\times$ subsampling in the 3D space and decode 4 Gaussians from each subsampled 3D point in our initial reconstruction model. Thus the number of Gaussians remains the same as previous per-pixel methods. The recurrent process remains the same as the many-view setups. Table 3 shows that our ReSplat outperforms previous feed-forward 3DGS models (e.g., DepthSplat (Xu et al., 2025), GS-LRM (Zhang et al., 2024a) and Long-LRM (Chen et al., 2025b)) by clear margins. We present visual comparisons in Fig.˜5, where our ReSplat produces better structures than MVSplat and DepthSplat. Compared to the 3DGS-free method LVSM (Jin et al., 2025), we outperform its encoder-decoder architecture by 1.1dB PSNR and our results are similar to its best-performing decoder-only model variant. However, our method offers the benefits of an explicit 3D Gaussian representation, enabling $20\times$ faster rendering speed.
Compression PSNR ↑ SSIM ↑ LPIPS ↓ Time (s)
@@
Table 4: Different compression factors. $16\times$ compression represents a good speed-accuracy trade-off.
Figure 6: Different compression factors.
Table 5: Ablations. We conduct several ablation experiments to validate our design choices.
@@
Full 27.83 0.880 0.122
w/o render error 26.77 0.865 0.142
w/o ResNet feature 27.78 0.879 0.128
w/o kNN attn 27.64 0.877 0.124
w/o global attn 27.77 0.879 0.122
@@ 4.2 Analysis and Ablation
Compression Factor. Be default, we compress the number of Gaussians by $16\times$ using depth maps at $1/4$ resolution. In Table˜4, we compare with $64\times$ (with $1/8$ depth maps) and $4\times$ (with $1/2$ depth maps) compression factors. We can observe that less compression leads to higher quality, as also visualized in Fig.˜6. However, $4\times$ compression is $2\times$ slower than $16\times$ for 8 input views at $256\times 448$ resolution. It would be more expensive when handling higher resolution images (e.g., $512\times 960$). Thus, we choose $16\times$ compression as a good speed-accuracy trade-off.
Ablation of the Initial Model. As shown in LABEL:tab:ablation_init, we observe that the kNN attention is crucial to maintain our performance when compressing the number of Gaussians by $16\times$. The global attention also brings moderate performance gains, which indicates that both local and global 3D context aggregation is important for learning compact 3D representations.
Ablation of the Recurrent Model. The rendering error plays an indispensable role in our recurrent model. As shown in LABEL:tab:ablation_refine, we didn’t observe any obvious improvement by removing the rendering error as input to our recurrent model. Computing the error in ResNet feature space is beneficial for the LPIPS metric, and both kNN attention and global attention contribute to the performance.
@@ 5 Conclusion
We present a recurrent Gaussian splatting model that achieves efficient and high-quality view synthesis in a feed-forward, gradient-free manner. By leveraging the rendering error as a feedback signal and operating in a compact subsampled 3D space, our method significantly reduces the number of Gaussians while improving performance and generalization across diverse scenarios.
@@
Limitations. Our current model relies on the kNN-based point attention (Zhao et al., 2021), which incurs high computational cost when the number of Gaussians becomes large (e.g., > 500K). We anticipate that more efficient point-based attentions (Wu et al., 2022; 2024) and sparse structures (Ren et al., 2024) could further improve the scalability and efficiency of our approach. In addition, our current model saturates after three iterations, and we speculate that keeping the number of Gaussians fixed during the recurrent process to be a potential reason and it would be interesting to design more adaptive update strategies and explore how to further scale up test-time compute. We view these directions as promising avenues for future research.
Acknowledgments. We thank Naama Pearl, Xudong Jiang, and Stefano Esposito for the insightful comments, and Yung-Hsu Yang and Kashyap Chitta for the fruitful discussions. Andreas Geiger was supported by the ERC Starting Grant LEGO-3D (850533) and the DFG EXC number 2064/1 - project number 390727645. This work was supported as part of the Swiss AI Initiative by a grant from the Swiss National Supercomputing Centre (CSCS) under project ID a144 on Alps.
Appendix A Additional Visualizations

### v2

@@ 4.1 Main Results
Table 1: Evaluation of 8 input views ($512\times 960$) on DL3DV. The standard optimization-based approach 3DGS (Kerbl et al., 2023) requires several thousands of iterations to reach convergence, while our feed-forward ReSplat is significantly faster and is able to benefit from additional iterations. Previous per-pixel feed-forward models MVSplat (Chen et al., 2024b) and DepthSplat (Xu et al., 2025) produces millions of Gaussians, while our ReSplat compresses the number of Gaussians by $16\times$, also leading to $4\times$ faster rendering speed.
@@
1 27.15 0.859 0.169 246K 0.437 0.0007
2 27.51 0.865 0.163 246K 0.563 0.0007
3 27.65 0.867 0.161 246K 0.789 0.0007
4 27.70 0.868 0.160 246K 0.816 0.0007
Figure 3: View synthesis on DL3DV. Our ReSplat outperforms both optimization and feed-forward methods, with significantly smaller rendering errors. More samples are presented in Fig.˜8 (appendix).
@@
8 Views at $512\times 960$ Resolution on DL3DV. We report the results on the DL3DV benchmark split (140 scenes) in Table˜1. For experiments with 3DGS (Kerbl et al., 2023), we perform per-scene optimization on the 8 input views for all the 140 scenes, while for feed-forward models, we perform zero-shot feed-forward inference. We observe that 3DGS optimization typically converges with 4K optimization steps, and optimizing longer could lead to overfitting due to the sparse input views, thus we report the best results at 4K iterations. As shown in Table˜1, 3DGS optimization is considerably slow due to the large number of iterations required, while our feed-forward ReSplat is $100\times$ faster and is able to benefit from additional iterations. Previous per-pixel feed-forward models MVSplat (Chen et al., 2024b) and DepthSplat (Xu et al., 2025) produces millions of Gaussians, while our ReSplat compresses the number of Gaussians by $16\times$, also leading to $4\times$ faster rendering speed. Overall, our ReSplat outperforms 3DGS by 4.2dB PSNR and DepthSplat by 3.5dB PSNR with superior efficiency on the number of Gaussians and the rendering speed. The visual comparisons are shown in Fig.˜3 and Fig.˜8, which demonstrate the higher rendering quality of our method. We provide additional visualizations of reconstructed Gaussians in Fig.˜12 and Fig.˜13 (appendix).
Optimization-Based vs. Feed-Forward Refinement. To further demonstrate the superiority of our feed-forward refinement, we compare with optimization-based 3DGS by using the same ReSplat’s initialization. From Fig.˜4 and Table˜8 (appendix), we can observe that our ReSplat is significantly faster than 3DGS optimization-based refinement thanks to our gradient-free and feed-forward nature.
(a) PSNR vs. number of iterations.
(b) Reconstruction time vs. number of iterations.
Figure 4: Optimization-based vs. feed-forward refinement. Starting from the same ReSplat’s initialization, we compare 3DGS optimization-based refinement with our feed-forward approach. Our ReSplat improves the rendering quality significantly faster (4 vs. 80 iterations) and is $53\times$ faster in terms of the reconstruction speed. See Table˜8 (appendix) for detailed numbers.
(a) Cross-dataset generalization.
(b) Cross-view generalization.
(c) Cross-resolution generalization.
@@
Figure 5: Generalization to unseen datasets, number of input views, and image resolutions. Our recurrent model demonstrates robust generalization despite being trained exclusively on DL3DV at a fixed resolution ($512\times 960$) with 8 input views. We note that the RealEstate10K dataset is easier than DL3DV due to smaller camera motion.
Generalization Across Datasets, View Counts, and Image Resolutions. We further evaluate the generalization abilities of our feed-forward model trained on exclusively DL3DV at $512\times 960$ resolution with 8 input views. First, when generalizing to the unseen RealEstate10K dataset (Fig.˜5(a)), the improvement brought by our recurrent model is more significant since unlike single-step feed-forward models (iteration 0), our model is able to adapt to the test data using the rendering error and thus reduces the domain gap. Second, we test our initial and recurrent models with more than 8 input views in Fig.˜5(b), and observe that our recurrent model benefits more from the additional input views, while the initial model tends to be saturated. This indicates that our rendering error-informed recurrent model exploits the additional information more effectively. Third, existing single-step feed-forward models usually suffer from obvious performance drop when the testing image resolution is different from training. However, with our recurrent model, we can significantly improve the robustness to different testing resolutions (Fig.˜5(c)). For example, our recurrent model improves 5dB PSNR when generalizing from $512\times 960$ to $320\times 640$. These experiments demonstrate the benefits of our recurrent model, which is able to adapt to the unseen test scenario using the rendering error as a feedback signal and thus substantially improves the robustness.
@@
Table 2: Evaluation of 16 input views ($540\times 960$) on DL3DV. Our ReSplat reconstructs $4\times$ fewer Gaussians than Long-LRM but still outperforms it.
@@
ReSplat 0 22.69 0.742 0.307 0.7sec 518K
1 23.23 0.758 0.291 1.2sec 518K
2 23.51 0.766 0.284 1.7sec 518K
16 Views at $540\times 960$ Resolution on DL3DV. We follow Long-LRM (Chen et al., 2025b) for this evaluation setup such that a direct comparison is possible. The results of 3DGS (Kerbl et al., 2023), Mip-Splatting (Yu et al., 2024), and Scaffold-GS (Lu et al., 2024) are borrowed from Long-LRM paper. This experiment aims to reconstruct the full DL3DV scene from 16 input views, which is particularly challenging due to the wide coverage of the DL3DV dataset. However, our ReSplat still outperforms previous optimization and feed-forward methods, as shown in Table˜2. Notably, Long-LRM uses Gaussian pruning based on the opacity values during training and evaluation, which leads to $\sim 4\times$ reduction of number of Gaussians. In contract, we compress the Gaussians by $16\times$, thus our final reconstruction has $4\times$ fewer Gaussians than Long-LRM while still outperforming it. Our reconstruction time is slower than Long-LRM mainly because of the kNN operation, which is expensive to compute with large number (e.g., >500>500K) of points. Further optimizing the implementation could potentially improve our reconstruction speed.
@@
ReSplat (Ours) 29.75 0.912 0.100
Table 3: Evaluation of two input views ($256\times 256$) on RealEstate10K. ReSplat outperforms prior feed-forward 3DGS models and matches LVSM’s quality with $20\times$ faster rendering.
@@
Table 4: RealEstate10K to ACID generalization.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
pixelSplat 27.64 0.830 0.160
MVSplat 28.15 0.841 0.147
DepthSplat 28.37 0.847 0.141
GS-LRM 28.84 0.849 0.146
ReSplat 29.87 0.864 0.135
Table 5: Different compression factors. $16\times$ compression represents a good speed-accuracy trade-off. The visual results are shown in Fig.˜7.
Compression PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time (s)
@@
2 Views at $256\times 256$ Resolution on RealEstate10K and ACID. Considering the redundancy is not a major issue with two input views at low resolution ($256\times 256$), we perform $4\times$ subsampling in the 3D space and decode 4 Gaussians from each subsampled 3D point in our initial reconstruction model. Thus the number of Gaussians remains the same as previous per-pixel methods. The recurrent process remains the same as the many-view setups. Table 3 shows that our ReSplat outperforms previous feed-forward 3DGS models (e.g., DepthSplat (Xu et al., 2025), GS-LRM (Zhang et al., 2024a) and Long-LRM (Chen et al., 2025b)) by clear margins. Compared to the 3DGS-free method LVSM (Jin et al., 2025), we outperform its encoder-decoder architecture by 1.1dB PSNR and our results are similar to its best-performing decoder-only model variant. However, our method offers the benefits of an explicit 3D Gaussian representation, enabling $20\times$ faster rendering speed. We present visual comparisons in Fig.˜6, where our ReSplat produces better structures than MVSplat and DepthSplat. In Table˜4, we show the zero-shot generalization results on the unseen ACID dataset, where our ReSplat again outperforms previous methods by clear margins.
Table 6: Analysis and Ablation. We conduct several experiments to validate our design choices.
Method PSNR ↑ SSIM ↑ LPIPS ↓
Initialization 26.77 0.865 0.142
w/o rendering error 27.19 0.873 0.137
RGB error only 27.90 0.882 0.130
Feature error only 28.77 0.897 0.110
Concat (RGB & feature errors) 28.93 0.900 0.106
Add (RGB & feature errors) 29.07 0.902 0.105
Reference PSNR ↑ SSIM ↑ LPIPS ↓
Initialization 26.77 0.865 0.142
COLMAP 28.14 0.886 0.116
First view 28.66 0.896 0.109
Last view 28.59 0.895 0.110
Middle view 29.07 0.902 0.105
@@
Full 29.07 0.902 0.105
w/o state 27.79 0.878 0.125
w/o kNN attn 28.58 0.894 0.111
w/o global attn 28.96 0.900 0.107
@@ 4.2 Analysis and Ablation
Rendering Error. We evaluate the effect of the rendering error in LABEL:tab:ablation_render_error. Removing the rendering error in our recurrent model leads to a significant performance drop (-1.9dB PSNR). The feature-space error performs better than the pixel-space RGB error, and the best performance is obtained by combining the RGB-space and feature-space errors (addition is slightly better than concatenation).
Coordinate System. We observe that aligning to the middle input view’s camera pose performs significantly better (+0.9dB PSNR) than using the default COLMAP’s global coordinate system. We attribute this to the temporal nature of video data, where anchoring to the middle frame acts as a central pivot, balancing the spatial distributions of the 3D Gaussians and making the learning task easier for the 3D network.
Compression Factor. Be default, we compress the number of Gaussians by $16\times$ using depth maps at $1/4$ resolution. In Table˜5, we compare with $64\times$ (with $1/8$ depth maps) and $4\times$ (with $1/2$ depth maps) compression factors. We can observe that less compression leads to higher quality, as also visualized in Fig.˜7 (appendix). However, $4\times$ compression is $2\times$ slower than $16\times$ for 8 input views at $256\times 448$ resolution. It would be more expensive when handling higher resolution images (e.g., $512\times 960$). Thus, we choose $16\times$ compression as a good speed-accuracy trade-off.
Ablation of the Initial Model. As shown in LABEL:tab:ablation_init, we observe that the kNN attention is crucial to maintain our performance when compressing the number of Gaussians by $16\times$. The global attention also brings moderate performance gains, which indicates that both local and global 3D contexts are important for learning compact 3D representations. The visual results are in Fig.˜10 (appendix).
Ablation of the Recurrent Model. The state input is pivotal to our recurrent network. Unlike the low-level, raw Gaussian parameters, it encodes high-level image and 3D features derived from our initialization model. Both kNN attention and global attention contribute to the performance. The visual results are in Fig.˜11 (appendix).
@@ 5 Conclusion
We have presented a feed-forward recurrent Gaussian splatting model that achieves efficient and high-quality view synthesis. By using the rendering error as an informative feedback signal and operating in a compact subsampled 3D space, our method significantly reduces the number of Gaussians while improving performance and generalization across datasets, view counts, and image resolutions.
@@
Limitations. Our current model relies on kNN-based point attention (Zhao et al., 2021), which incurs high computational cost when the number of Gaussians becomes large (e.g., > 500K). We anticipate that more efficient point-based attentions (Wu et al., 2022; 2024) and sparse structures (Ren et al., 2024) could further improve the scalability and efficiency of our approach. In addition, our current model saturates after four iterations, it would be interesting to explore how to further scale up test-time compute. We view these directions as promising avenues for future research.
Appendix A Additional Evaluations
Comparison with 3DGS Across 8, 16, and 32 Views. In Table˜7, we compare with optimization-based 3DGS across 8, 16 and 32 input views. The quality gap becomes smaller when given 32 views for optimization-based 3DGS. However, our ReSplat is more than $200\times$ faster in terms of the reconstruction speed. In this setup, we expand the sampling region as the number of input views increases to enlarge scene coverage. Consequently, the test views differ across configurations to account for this larger spatial extent.
Optimization-Based vs. Feed-Forward Refinement. In Table˜8, we show the detailed numbers of comparing 3DGS optimization-based and our feed-forward refinement starting from the same ReSplat’s initialization. See Fig.˜4 (main paper) for the visualization of this table.
Recurrent vs. Non-recurrent Architecture. We demonstrate the effectiveness of our recurrent model by comparing with non-recurrent variants in Table˜9. In particular, we first compare with non-weight-sharing multi-step stacked networks where different iterations have different model weights and all the other components are the same. We can observe that non-weight-sharing not only leads to $4\times$ more parameters, but also results in worse view synthesis results. We further compare with non-weight-sharing single-step deeper networks by increasing the number of attention blocks for a single-step refinement, where the results are clearly worse than our multi-step recurrent network. The weight-sharing design in our recurrent network implicitly regularizes training, which is not only more parameter-efficient but also leads to better results.
Comparison with SplatFormer. We note that our ReSplat has several crucial differences with SplatFormer (Chen et al., 2025a). First, we identify the rendering error as an informative feedback signal for improving the Gaussian splats, which is missing in SplatFormer. Second, ReSplat is a recurrent model which supports multi-step refinement with a weight-sharing architecture, while SplatFormer is a non-recurrent network designed for single-step refinement. Third, ReSplat is a pure feed-forward model with feed-forward initialization and feed-forward refinement. In contrast, SplatFormer relies on lengthy optimization-based 3DGS to get initial Gaussians. Fourth, SplatFormer is designed for object-centric datasets, while ReSplat can handle diverse scene-level datasets where the complexity is much higher than objects. Despite these differences, we tried to conduct a comparison with SplatFormer on our scene-level datasets. We found it particularly challenging to make it work properly for scene-level datasets, since it replies on Point Transformer V3 (Wu et al., 2024) where a proper grid size is required to serialize the point cloud. This can be done for object-centric datasets where normalizing the objects to [1,1][-1,1] is possible. However, for unbounded scene-level datasets, this would be very challenging. We tried different normalizations and did grid search for the grid size, and the best results we obtained with SplatFormer are reported in Table˜10. We can see that SplatFormer is 2dB PSNR worse than our method.
Comparison with Sparse View Optimization Methods. We additionally compare with optimization-based 3DGS methods that are specifically designed for sparse input views. These sparse-view optimization methods (Li et al., 2024; Chung et al., 2024) usually rely on additional depth losses to regularize the optimization process. To compare with them, we perform 3DGS optimization with an additional depth loss between the rendered depth map and the estimated monocular depth map from Depth Anything V2 (Large) (Yang et al., 2024). The results are reported in Table˜11. With the additional depth loss, the 3DGS optimization results are improved by 1dB PSNR. However, the gap with our ReSplat is still significant (3dB PSNR). The additionally introduced depth loss also makes the optimization slower due to the additional time for depth rendering and monocular depth estimation. In contrast, our model doesn’t rely on any additional supervision from an external monocular depth model and it’s $94\times$ faster thanks to our feed-forward nature.
Features for Computing the Rendering Error. In Table˜12, we compare ResNet features with those from DINOv2. We observed no improvement when using the larger, more recent feature extractor. We attribute this to the patch-based architecture of DINOv2, which may result in coarser spatial information. In contrast, convolutional networks maintain local structural fidelity, which is critical for high-quality pixel-accurate view synthesis.
Model Profiling. In Table˜13, we report the total runtime and individual component latency. In the initial reconstruction model, the depth prediction module constitutes the majority of the runtime. For the recurrent model, the kNN attention mechanism consumes the most time. These results highlight potential areas for future optimization.
Table 7: Comparison with optimization-based 3DGS across 8, 16, and 32 input views. The quality gap becomes smaller when given 32 views for optimization-based 3DGS. However, our ReSplat is more than $200\times$ faster in terms of the reconstruction speed. The image resolution is $256\times 448$.
#Views Method Category #Iterations PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ #Gaussians
Recon.
Time (s)
8 3DGS Optimization 4000 26.44 0.841 0.134 250K 49
ReSplat Feed-Forward 4 29.20 0.904 0.104 57K 0.21
16 3DGS Optimization 4000 27.38 0.864 0.119 395K 70
ReSplat Feed-Forward 4 29.01 0.900 0.105 114K 0.34
32 3DGS Optimization 4000 27.86 0.879 0.113 522K 160
ReSplat Feed-Forward 4 28.30 0.891 0.114 229K 0.75
Table 8: Optimization-based vs. feed-forward refinement. Starting from the same ReSplat’s initialization, we compare 3DGS optimization-based refinement with our feed-forward approach. Our ReSplat improves the rendering quality significantly faster (4 vs. 80 iterations) and is $53\times$ faster in terms of the reconstruction speed. See Fig.˜4 (main paper) for the visualization of this table.
Method # Iter. PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time (s) $\downarrow$
ReSplat (init) 0 26.21 0.842 0.185 0.311
ReSplat (init) + 3DGS 1 26.29 0.844 0.185 0.852
ReSplat (init) + 3DGS 2 26.44 0.846 0.183 1.393
ReSplat (init) + 3DGS 3 26.55 0.848 0.183 1.934
ReSplat (init) + 3DGS 4 26.64 0.849 0.183 2.475
ReSplat (init) + 3DGS 10 26.98 0.855 0.183 5.721
ReSplat (init) + 3DGS 20 27.29 0.960 0.183 11.131
ReSplat (init) + 3DGS 40 27.56 0.863 0.181 21.951
ReSplat (init) + 3DGS 60 27.68 0.865 0.180 32.771
ReSplat (init) + 3DGS 80 27.73 0.865 0.179 43.591
ReSplat 1 27.15 0.859 0.169 0.439
ReSplat 2 27.51 0.865 0.163 0.567
ReSplat 3 27.65 0.867 0.161 0.795
ReSplat 4 27.70 0.868 0.160 0.823
Appendix B Additional Details
Training Details. We train our model with cosine learning rate schedule. For experiments on DL3DV, we adopt a progressive training strategy with gradually increased image resolutions and number of input views for better efficiency. More specifically, we first train our model with 8 input views at $256\times 448$ resolution, and then we fine-tune the model with 8 input views at $512\times 960$ resolution, and finally we fine-tune the model with 16 input views at $512\times 960$ resolution. For each stage, we train with 16 GH200 GPUs for 80K steps, with 50K steps for the initial reconstruction model and 30K steps for the recurrent model. For experiments on RealEstate10K at $256\times 256$ resolution, we first train the initial model with 16 GH200 GPUs for 200K steps and then train the recurrent model for 100K steps.
Appendix C Additional Visualizations
We invite readers to our supplementary video to view our video rendering results.
In Fig.˜8, we show more visual comparisons with 3DGS, MVSplat and DepthSplat on DL3DV dataset.
In Fig.˜9, we show the visual results with different numbers of iterations.
In Fig.˜10, we show the visual ablation results of our initial model. In Fig.˜11, we show the visual ablation results of our recurrent model.
In Fig.˜7, we show the visual results of different compression factors, and the detailed metrics are in Table˜5 (main paper). We choose $16\times$ compression since it offers a good trade-off between speed and accuracy.